A particle moves along the x axis according to the equation x ! 2.00 % 3.00t # 1.00t2, where x is in meters and t is in seconds. At t ! 3.00 s, find (a) the position of the particle, (b) its velocity, and (c) its acceleration.
1
Expert's answer
2020-11-23T10:32:28-0500
"x=2+3t-t^2"

a)


"x(3)=2+3(3)-3^2=2\ m"

b)


"v(3)=3-2(3)=-3\frac{m}{s}"

c)


"a(3)=-2\frac{m}{s^2}"